# Improving Low-Resource Machine Translation via Round-Trip Reinforcement Learning

### v1

@@ Abstract
Low-resource machine translation (MT) has gained increasing attention as parallel data from low-resource language communities is collected, but many potential methods for improving low-resource MT remain unexplored. We investigate a self-supervised reinforcement-learning-based fine-tuning for translation in low-resource settings using round-trip bootstrapping with the No Language Left Behind (NLLB) family of models. Our approach translates English into a target low-resource language and then back into English, using a combination of chrF++ and BLEU as the reward function on the reconstructed English sentences. Using the NLLB-MD dataset, we evaluate both the 600M and 1.3B parameter NLLB models and observe consistent improvements for the following languages: Central Aymara, Friulian, Wolof and Russian. Qualitative inspection of translation outputs indicates increased fluency and semantic fidelity. We argue that our method can further benefit from scale, enabling models to increasingly leverage their pretrained knowledge and continue self-improving.
@@ 4 Experiments
We base our experiments on NLLB-200 (No Language Left Behind) Goyal and others (2022), a massively multilingual translation model that covers 200 languages. In particular, we use the distilled 600M and 1.3B parameter variants. NLLB is an encoder-decoder Transformer Vaswani and others (2017) with language-specific tokens that specify the target language during generation. We use the NLLB-MD dataset Guzman et al. (2019); Goyal and others (2022) in all our experiments. We use the provided split of parallel train-, dev- and test-set with respective sizes of 6000, 1310 and 1600 sentences. More on hyperparameters and implementation details are in appendix A. Our study focuses on four languages: Central Aymara, Friulian, Wolof and Russian. We use GoldFish Chang et al. (2024) monolingual model family to score the fluency of the forward translations for every language in addition to chrF++.
First, we compare the base and trained models on these four languages using chrF++ and additionally translation fluency using Goldfish log-probabilities Chang et al. (2024). Next, we present reward trajectories for forward and backward translation directions. Also, we present a reward function abalation study. Finally, we provide a qualitative analysis of model outputs on examples that are not seen in the training data.
@@ 5 Results
600M 1.3B
Language Before After Before After
Central Aymara 24.38 27.96 (+3.58) 26.01 28.70 (+2.69)
Friulian 45.75 47.50 (+1.75) 49.41 50.74 (+1.33)
Wolof 20.73 23.78 (+3.05) 24.73 26.81 (+2.08)
Russian 48.80 52.68 (+3.88) 50.10 54.21 (+4.11)
Table 1: Forward Translation chrF++ scores on the test set before and after training on each language for the 600M and 1.3B models using the composite reward function (Bleu + chrF++). Green values indicate absolute improvement in parentheses.
@@ 5.1 Forward Translations Evaluation
The largest gain is observed on Russian (+3.88), suggesting that the intrinsic round-trip signal strongly improves adequacy and consistency even when the target language is comparatively well supported in multilingual pretraining. It’s also evident that the more knowledge the model has on the language the better it is able to exploit it with the chrF++ reward signal.
For the 1.3B model, we observe gains of +1.33 on Friulian, +2.08 on Wolof, +4.11 on Russian, and +2.69 on Central Aymara. Notably, the largest improvement again occurs on Russian (+4.11), indicating that increased model capacity does not saturate the benefits of our self-consistency objective; instead, larger models appear to better exploit the reward signal.
@@ 5.5 Qualitative Analysis On Translation Outputs
Finally, in Table 4 we provide a qualitative comparison of translations produced by the 1.3B base model and the trained model across four evaluation languages (one example each; additional samples are reported in Appendix B). Overall, the trained model produces outputs that are close to the human reference.
@@ 8 Ethical Considerations
Appendix A Implementation Details
We implement our method using PyTorch Paszke et al. (2019) and the Hugging Face Transformers Wolf et al. (2020) library. Table 5 summarizes the key hyperparameters used in our experiments.
We use AdamW Kingma and Ba (2017) for optimization. During generation, we employ nucleus sampling with temperature scaling to encourage diverse candidate generation.
Hyperparameter Value
Learning rate $2\times 10^{-6}$
Batch size 2
Group size $K$ 4
Generation temperature 1.8
Top-$k$ sampling 100
Top-$p$ sampling 0.95
KL coefficient $\beta$ 0.04
Clip parameter $\epsilon_{c}$ 0.2
Reference update frequency 16 steps
Maximum sequence length 64 tokens
Training epochs 2
Table 5: Hyperparameters used for training.
Appendix B Qualitative Translation Examples
B.1 Central Aymara
@@ B.4 Wolof
Appendix C Additional Training Plots

### v2

@@ Abstract
Low-resource machine translation (MT) has gained increasing attention as parallel data from low-resource language communities is collected, but many potential methods for improving low-resource MT remain unexplored. We investigate a self-supervised reinforcement-learning-based fine-tuning for translation in low-resource settings using round-trip bootstrapping with the No Language Left Behind (NLLB) family of models. Our approach translates English into a target low-resource language and then back into English, using a combination of chrF++ and BLEU as the reward function on the reconstructed English sentences. Using the NLLB-MD dataset, we evaluate both the 600M and 1.3B parameter NLLB models and observe consistent improvements for the following languages: Central Aymara, Friulian, Wolof and Russian. Qualitative inspection of translation outputs indicates increased fluency and semantic fidelity. We argue that our method can further benefit from scale, enabling models to increasingly leverage their pretrained knowledge and continue self-improving. The code is available on github: https://github.com/Copticoder/thesis-nllb-bootstrap-grpo.
@@ 3.3 Reward Design
Table 1: Main results (higher is better). For each language and model size, we bold the best score and underline the second-best across {Base, BT, UMT, Ours}. BT denotes back-translation, UMT denotes unsupervised MT using only monolingual corpora, and Ours denotes round-trip MT with GRPO.
NLLB-600M NLLB-1.3B
Language Base BT UMT Ours Base BT UMT Ours
Central Aymara 24.38 27.34 24.85 28.13 26.01 28.61 28.21 28.73
Wolof 20.73 23.57 19.89 23.96 24.73 26.35 27.72 26.66
Russian 48.80 53.31 51.70 52.79 50.10 55.46 55.14 55.16
Friulian 45.75 47.38 48.96 47.61 49.41 51.72 46.43 50.76
Avg. 34.91 37.90 36.35 38.12 37.56 40.53 39.38 40.33
@@ 4 Experiments
4.1 Our Setting
We base our experiments on NLLB-200 (No Language Left Behind) Goyal and others (2022), a massively multilingual translation model that covers 200 languages. In particular, we use the distilled 600M and 1.3B parameter variants. NLLB is an encoder-decoder Transformer Vaswani and others (2017) with language-specific tokens that specify the target language during generation. We use the NLLB-MD dataset Guzman et al. (2019); Goyal and others (2022) in all our experiments, and follow the provided split of parallel train, dev, and test sets with respective sizes of 6000, 1310, and 1600 sentences. We implement our method in PyTorch Paszke et al. (2019) using the Hugging Face Transformers library Wolf et al. (2020). We optimize with AdamW Kingma and Ba (2017) using a learning rate of $2\times 10^{-6}$ for 2 epochs with batch size 2, maximum sequence length 64, group size K=4K{=}4, KL coefficient $\beta{=}0.04$, clip parameter $\epsilon_{c}{=}0.2$, and a reference update frequency of 16 steps (Table LABEL:tab:hyperparams). During generation, we use nucleus sampling with temperature scaling (temperature 1.8, top-$k$ 100, top-$p$ 0.95) to encourage diverse candidate generation. Our study focuses on four languages: Central Aymara, Friulian, Wolof, and Russian. We use the GoldFish Chang et al. (2024) monolingual model family to score the fluency of the forward translations for each language in addition to chrF++.
4.2 Baselines
Unsupervised Machine Translation Using Monolingual Corpora Only (round-trip / cycle training).
We include Lample et al. (2018) method with denoising and cycle-consistent reconstruction objectives, augmented with an adversarial discriminator. Models are trained for $2$ epochs with batch size $2$. We use separate learning rates for the encoder/decoder and discriminator: $\text{LR}_{\text{enc/dec}}=3\cdot 10^{-5}$ and $\text{LR}_{\text{disc}}=5\cdot 10^{-5}$. Objective weights are set to $\lambda_{\text{auto}}=1.0$ (denoising auto-encoding), $\lambda_{\text{cd}}=1.0$ (cycle/round-trip reconstruction), and $\lambda_{\text{adv}}=1.0$ (adversarial alignment). Input corruption uses word dropout $0.1$ and local shuffle with window size $k=3$.
Back-Translation
We use normal back-translation Sennrich et al. (2016) with supervised learning on the dummy back-translated sentences. Moreover, we translate from the source English into the dummy low-resource language translations, and we use this to back-translate into English and do normal cross entropy loss on the original and back-translated English text. We use back-translation batch size of 2 and $\text{LR}_{\text{disc}}=3\cdot 10^{-5}$.
@@ 5.1 Forward Translations Evaluation
Our method is particularly effective for very low-resource languages. With the 600m model, it achieves the best scores on both Central Aymara (28.13) and Wolof (23.96), outperforming BT by 0.79 and 0.39 points respectively, and substantially exceeding UMT. With the 1.3B model, our method remains best on Central Aymara (28.73) and is competitive on Wolof (26.66), where UMT attains the top score (27.72). For higher-resource languages (Russian, Friulian), BT is often the strongest baseline, but the gap to our method is small (e.g., within 0.52 on Russian at 600M and within 0.30 at 1.3B). On average, our method is best for the 600m model (38.12), and is within 0.20 of the best method for for the 1.3B model (40.33 vs. 40.53 for BT). Finally, UMT is less consistent across model sizes and languages, which is plausibly explained by its more complex optimization objective that combines auto-encoding, cross-domain, and adversarial losses.
@@ 5.5 Qualitative Analysis On Translation Outputs
Finally, in Table 4 we provide a qualitative comparison of translations produced by the 1.3B base model and the trained model across four evaluation languages (one example each; additional samples are reported in Appendix A). Overall, the trained model produces outputs that are close to the human reference.
@@ 8 Ethical Considerations
Appendix A Qualitative Translation Examples
A.1 Central Aymara
@@ A.4 Wolof
Appendix B Additional Training Plots